# Interpreting the Linear Structure of Vision-language Model Embedding Spaces

### v2

@@ 6 Analysis 2: Are We Finding Features With Cross-Modal Meaning?
Most Concepts Are Single-Modality
Figure 4: Histogram of the modality scores (top) and the modality scores weighted by energy (bottom) of every concept in the SAEs that we trained for all four models. On every model, the modes are the two extremes: concepts either activated only by text or only by image. However, weighting by energy makes this much more prominent, showcasing that almost all of the reconstruction coefficients go to concepts that are single-modality.
@@
Cross-Modal Alignment from Unimodal Bases
Interestingly, the predominance of single-modality concepts in the SAE dictionary does not preclude the model’s ability to align across modalities. This sets up a central tension: if the concepts that dominate reconstruction are modality-specific, how does the model support semantic connections between paired image–text inputs? The resolution lies in the structure of the SAE dictionary itself. While most concepts activate primarily on one modality, they do not exist in isolation. The sparse autoencoder enables bridging via geometrically aligned and co-activated concept pairs—concepts that live in distinct modality clusters but consistently respond to semantically related inputs.
Finding: Despite being functionally unimodal, concepts can form semantically meaningful cross-modal bridges through high co-activation and directional alignment. These bridges support joint understanding across text and image even without shared concepts.
As illustrated in Figure 5, high-bridge-score connections often capture many-to-many relationships—for example, a visual concept like “surfboards” aligns not just to a single caption, but to a diverse cluster of related descriptions. This illustrates how the model achieves flexible semantic grounding across modalities using sparse but structured interconnections.
Many Concepts Are Nearly Orthogonal to the Modality Subspace
Given that most concepts are single-modality in usage, one might ask whether they are also geometrically aligned with modality – i.e., whether they lie along directions that distinguish image from text embeddings. If so, this would suggest that the concepts primarily reflect modality separation, rather than shared or cross-modal meaning. We begin by analyzing the embedding space itself. Prior work has shown that representations in deep models often occupy anisotropic regions – narrow cones within the full space (Ethayarajh, 2019; Mimno & Thompson, 2017). In our setting, we ask whether image and text embeddings occupy separate cones (Tyshchuk et al., 2023): we compute the average cosine distance between pairs of embeddings from the same modality versus different modalities. As shown in Figure 6, embeddings of the same modality (green, blue) are significantly closer than those of different modalities (orange). This suggests that image and text activations reside in distinct conical regions of the embedding space.
@@
Figure 6: Within-modality anisotropy for CLIP and SigLIP. Cosine similarities show that image–image and text–text pairs (blue/green) are closer than image–text pairs (orange), indicating that embeddings occupy separate modality-specific cones.
Do the concept vectors align with the modality structure? If they follow the geometry of the modality cones—separating image from text—they likely encode little cross-modal content. But if many lie near directions orthogonal to the modality subspace, they may instead reflect shared meaning. We test this by using each concept as a linear modality classifier: we project activations onto the concept direction and measure how well it separates image from text. High accuracy implies alignment with modality; near-chance performance suggests the concept is modality-agnostic and may encode cross-modal structure.
@@
Figure 7: Histogram of the accuracies of concepts used as classifiers, with the left side weighting each concept in the histogram by its energy. An accuracy of 0.5 means that the concept is orthogonal to every linear direction that defines modality (and therefore likely encodes some aspect of meaning independent from modality), and an accuracy of 1 means that the concept is aligned with a modality direction. We see that, especially in CLIP and AIMv2, there is a significant proportion of concepts that are almost orthogonal to the modality subspace. When we weight by energy, we see that accuracy-1 concepts are generally very high-energy across models, but that concepts less aligned with modality still receive a significant amount of energy.
@@
Figure 7 presents the distribution of classification accuracies for all concepts (left) and weighted by energy (right). While many concepts—particularly high-energy ones – are highly predictive of modality, a substantial number achieve near-random accuracy. These findings suggest that although most energetic concepts are modality-specific in activation, many are geometrically close to a modality-agnostic subspace. This points to the existence of sparse linear directions that support semantic alignment across modalities while remaining sensitive to modality itself.
Figure 8: The SAE projection effect. Even if a concept direction is nearly orthogonal to modality, the SAE encoder projects activations through $\bm{Z}=\bm{\Pi}\left\{\bm{A}\bm{W}+\bm{b}\right\}$, selecting the top-scoring concepts per input. Because the input distributions differ by modality, the thresholding step (e.g., TopK) may consistently select the concept for one modality but not the other. This creates functionally single-modality concepts, even when their directions lie near the modality-agnostic subspace.
How can concepts be both single-modality and almost orthogonal to modality? The SAE projection effect.
The results reveal an apparent contradiction: many concepts activate almost exclusively for one modality (Figure 4), yet their directions are poor modality classifiers (Figure 7), suggesting they lie near the modality-agnostic subspace. How can a concept be geometrically neutral yet functionally single-modality?
The resolution lies in the SAE projection step: activations are encoded as $\bm{Z}=\bm{\Pi}\left\{\bm{A}\bm{W}+\bm{b}\right\}$, where $\bm{\Pi}\left\{\cdot\right\}$ enforces sparsity (e.g., TopK or ReLU). Even if a concept direction is nearly orthogonal to modality, one modality may consistently produce higher pre-activation scores, causing it to be selected more often. This leads to modality-specific usage driven by modality-skewed thresholding, rather than geometry. We illustrate this effect in Figure 8.
Finding: A concept direction can be almost orthogonal to modality in direction, yet be functionally single-modality in activation, due to the interaction between input distribution and the sparsity projection.
@@ 7 Discussion and Conclusion
We study the linear structure of vision-language embedding spaces using SAEs trained on four VLMs. While concept dictionaries vary across seeds and data mixtures, the high-energy subset is consistently recovered and accounts for nearly all reconstruction. Most concepts are unimodal in usage but often lie near the modality-agnostic subspace, suggesting shared structure shaped by SAE thresholding rather than direction. To analyze cross-modal alignment, we introduce the Bridge Score, identifying concept pairs that are both geometrically aligned and co-activated. We release VLM-Explore, an interactive tool to visualize these structures across models.
@@ Appendix A Additional Figures and Expanded Discussion
We propose to supplements the main analyses by providing extended visualizations and detailed commentary for all four models—CLIP, SigLIP, SigLIP2, and AIMv2—used throughout the paper. Each subsection corresponds to a specific analysis presented in Section 5 and Section 6 of the main paper.
A.1 UMAP Visualizations Across Data Mixtures
@@
A.2 Modality Score Histograms
Figure 10: Modality score histograms for all models. Top: unweighted. Bottom: weighted by energy. Related to Figure 4.
The histograms in Figure 10 provide an empirical distribution over modality scores across all four models. The top row shows raw counts of modality scores without weighting; the bottom row weights each concept by its energy.
We observe two critical patterns:
  • The modality score distribution is clearly bimodal across all models: most concepts are either strongly image-specific (score near 1) or text-specific (score near 0). This confirms that although the joint embedding space is shared across modalities, its internal structure is dominated by modality-specific axes.
  • When weighted by energy, this bimodality becomes more pronounced: nearly all reconstruction-relevant concepts are unimodal. That is, high-energy concepts are not just modality-skewed : they are overwhelmingly unimodal in practice.
These results extend the findings of Section 6 and illustrate that, while multimodal alignment exists, it emerges from indirect interactions among strongly unimodal building blocks.
A.3 Within-Modality Anisotropy
Figure 11: Within-modality anisotropy: Cosine similarity distributions across image-image, text-text, and image-text embedding pairs. Related to Figure 6.
Figure 11 presents the cosine similarity distributions between pairs of embeddings from the same and different modalities for each model. These distributions reveal that image-image and text-text embedding pairs consistently exhibit higher mutual similarity than image-text pairs, supporting the conical anisotropy hypothesis discussed in the main paper.
This structure—text and image activations occupying separate cones—is consistent across CLIP, SigLIP, SigLIP2, and AIMv2. It confirms that the VLMs, despite being trained for modality alignment, still maintain a strongly separable embedding geometry.
A.4 Concept Classification Accuracy
Figure 12: Histogram of concept-wise modality classification accuracy for all models. Top: unweighted. Bottom: weighted by energy. Related to Figure 7.
Figure 12 shows the results of using each concept as a linear modality classifier—testing how well projection onto a concept vector can separate image vs. text embeddings.
Insights from this analysis:
  • Many concepts (particularly low-energy ones) perform at near-random accuracy, suggesting that they lie close to the modality-agnostic subspace and may encode more abstract or shared meaning.
  • High-energy concepts, however, often act as strong modality classifiers (accuracy near 1), especially in CLIP and AIMv2. Yet, importantly, not all high-energy concepts are modality-aligned indicating that semantically meaningful, shared features may still contribute to reconstruction.
These findings support the nuanced picture presented in Section 6: although usage is often unimodal, some concept vectors are not aligned with the modality separation direction, suggesting a functional dissociation between activation and geometric alignment.
@@ Appendix B Illustrative Examples from the Interactive Demo
We now present examples from the VLM-Explore interface (Section 4), illustrating concrete semantic content of discovered concepts and their inter-modal associations via bridge scores.
Figure 13: Interactive Demo: Example of text concept “banana” and associated visual concepts in CLIP, discovered using the bridge score. Shows clear cross-modal alignment despite being distant in UMAP space (non-trivial associations).
Figure 13 shows the concept cluster for the keyword “banana” as rendered in the VLM-Explore interface. Text instances grouped under this concept activate strongly for semantically aligned visual representations. The bridge score identifies connected visual concepts, which often correspond to similar object-level semantics (e.g., bananas on a table, being held, or cut). This demonstrates that the bridge score successfully identifies semantically aligned relations.
Figure 14: More aligned concept examples (CLIP): Object and scene-level text concepts and their linked visual activations. Shows concept-level semantic granularity.
These examples show how the bridge framework recovers granular semantic regularities across modality, spanning categories like scenes, humans, and objects. Despite functional single-modality usage (due to projection effects), the bridge allow to show the practical alignment.
Figure 15: Relational concepts: “cat sitting,” “man skateboarding,” and “holding umbrella.” These examples reflect the model’s ability to capture spatial or action-based semantics.
The bridge score also captures compositionality. In Figure 15, relational and action-based meanings are represented—text concepts describing activities map to corresponding image concepts, indicating joint semantic structure beyond raw modality.
Figure 16: Concepts corresponding to texture and color: “black and white” and “wooden.” These illustrate the fine-grained properties encoded by individual SAE features.
Finally, Figure 16 demonstrates the interpretability of concepts capturing non-object attributes – e.g., texture, material, and tone. These show that the SAE not only captures objects or scenes, but also abstract perceptual features.

### v3

@@ 5 Analysis 2: Are We Finding Features With Cross-Modal Meaning?
5.1 Most concepts are single-modality
@@
Figure 5: Image and text activations lie in separate cones. We plot the distribution of cosine similarities between model activations. Image–image and text–text pairs (blue/green) are close, while image-text pairs (orange) are much farther, indicating that embeddings occupy separate modality-specific cones.
5.2 Image and text activations lie in separate cones — but many concepts are orthogonal to this difference
@@
Figure 6: Many concepts are not aligned with the modality directions. Histogram of the accuracies of concepts used as classifiers (on the bottom each concept is weighted by its energy). An accuracy of 0.5 means that the concept is orthogonal to every linear direction that defines modality (and therefore likely encodes some aspect of meaning independent from modality), and an accuracy of 1 means that the concept is aligned with a modality direction. We see that, especially in CLIP and AIMv2, there is a significant proportion of concepts that are almost orthogonal to the modality subspace. When we weight by energy, we see that accuracy-1 concepts are high-energy, but that concepts less aligned with modality still receive a significant amount of energy.
In all models that we analyze, we find that the activations of image and text inputs lie in separate, narrow cones. To measure this, we can look at the cosine similarities between activations of image and text inputs, and how this compares to the cosine similiarities between activations of the same modality. If the average cosine similarity between a set of points is significantly different from zero, this means that the points lie in a narrow cone with respect to the origin, rather than being distributed around the unit ball evenly. Following findings by Tyshchuk et al. (2023) (as well as Mimno & Thompson (2017) and Ethayarajh (2019) in the language-only context), we find that image and text activations lie on separate, narrow cones. We present our results in Figure 5: embeddings of the same modality (green, blue) are significantly closer than those of different modalities (orange), which are on average almost orthogonal. This suggests that image and text activations reside in distinct conical regions of the embedding space.
@@
Do the concept vectors align with the modality structure? If they follow the geometry of the modality cones that separate image from text, they likely encode little cross-modal content. But if many lie near directions orthogonal to the modality subspace, they may instead reflect shared meaning. We test this by testing how well each concept performs as a linear modality classifier. To do this, we project input activations onto the direction defined by the concept, and measure how well it separates image from text. High accuracy implies alignment with modality; near-chance performance suggests the concept is modality-agnostic and may encode cross-modal structure. Figure 6 presents the distribution of classification accuracies for all concepts (left) and weighted by energy (right). While many concepts—particularly high-energy ones – are highly predictive of modality, a substantial number achieve near-random accuracy. These findings suggest that many concepts geometrically close to a modality-agnostic subspace, even though they activate for a specific modality This points to the existence of sparse linear directions that support semantic alignment across modalities while remaining sensitive to modality itself.
5.3 Proposal: Cross-modal alignment in VLM linear concepts
Our results lead us to a proposal of how the linear directions found by SAEs in multimodal space function, of which we provide a schematic in Figure 7
  1. 1.
    The image and text activations in the VLMs we study lie in separate, narrow cones (Figure 5), and SAE concepts tend to activate primarily for one modality (Figure 4).
  2. 2.
    However, SAE concepts largely do not lie along the linear subspace that separates the image and text cones (Figure 6), and in practice they can express semantics bridge between modalities. The Bridge Score (Equation 6) shows how there are many-to-many relationships between related concepts in text and image space (see Appendix B, Figure 16 for an illustrative example)
  3. 3.
    How can we resolve this tension? The resolution lies in considering the SAE projection effect (Figure 8): due to the sparsity constraints (like TopK), SAE concepts only activate for a tiny number of the input activations that they are well-aligned with. If an SAE concept is almost orthogonal to the modality-separating subspace, but leans slightly towards text, it is very unlikely to ever be in the top-K concepts for an image input. Though the concept will seem like a text-only concept, it is in fact aligned with many image outputs.
Figure 7: A schematic of our proposal in a 3D space. Though images and text can lie in separate orthogonal cones with respect to some subspace (in this case the x-y plane), cross-modal directions can emerge (in this case, the z direction coming towards us represents the cross-modal semantics of “giraffe”). In large latent spaces, many dimensions can form a complex multimodal space that the z-axis is standing in for here.
Figure 8: The SAE projection effect: We plot the dot product of a set of text (blue) and image (orange) activations with an SAE concept. Even though the concept direction does not separate modality, the TopK thresholding step (dotted line) may consistently select only one modality
6 VLM-Explore: An interactive, functional concept explorer
Lastly, we present VLM-Explore (https://vlm-concept-visualization.com), an interactive visualization tool designed to facilitate the exploration and analysis of our SAE concept representations. VLM-Explore offers an intuitive visualization that combines into one interactive figure four important aspects that help us interpret and debug VLMs. These are: 1) the UMAP structure shows how the linear directions of the latent space relate to each other 2) the maximally activating examples panel shows what each linear feature actually represents 3) Each feature is colored according to its Modality Score (Equation 5) showing how the two modalities share the space, and 4) the Bridge Score connections show what what semantics connect the two modalities.
@@ 7 Discussion and Conclusion
We study the linear structure of vision-language embedding spaces using SAEs trained on four VLMs. While concept dictionaries vary across seeds and data mixtures, the high-energy subset is consistently recovered and accounts for nearly all reconstruction. Most concepts are single-modality in usage but often lie near the modality-agnostic subspace, suggesting shared structure and a modality score shaped by SAE thresholding rather than direction. To analyze cross-modal alignment, we introduce the Bridge Score, identifying concept pairs that are both geometrically aligned and co-activated. We release VLM-Explore, an interactive tool to visualize these structures across models.
@@ Appendix A Additional Figures and Expanded Discussion
A.1 UMAP Visualizations Across Data Mixtures for all models
@@ Appendix B Illustrative Examples from the Interactive Demo
B.1 Examples of concepts
We present some qualitative examples of what kinds of concepts the SAEs that we train in the paper fire for, showing concepts from all 4 models. We showcase concepts of different abstractions, starting from more surface-level concepts and going to more abstract concepts. These concepts can all be accessed by putting their number into the Specific ID field in https://vlm-concept-visualization.com/ after selecting the correct VLM from the drop-down.
Figure 10: Coherent concept, with similar visual profile. Concept #2261 in SigLIP2 fires for pictures of pasta and broccoli, which is an interpretable concept, where all of the images look pretty similar in terms of color, texture, and shape composition
Figure 11: Coherent concept, with different visual profiles. Concept #5114 in AIMv2 fires for pictures of people getting haircuts. Though these pictures have fewer surface-level elements that are consistent between them, they all depict the same higher-level interpretable event.
Figure 12: Similar visual profiles, different semantics. Concept #1033 in CLIP fires for pictures that have vintage film coloring, rounded edges, and are outdoors. Though the surface-level visual profiles of the top activating examples for this concepts are related, there aren’t discernible higher-level semantics that connect them.
Figure 13: Higher-level abstract semantic concept. Concept #5812 in SigLIP fires for pictures of many disorganized things. The concepts which are most related to it by BridgeScore are text concepts detailing the idea of many things, rather than any other more surface-level semantics of the images (like, “toys”), hinting that the model has learned a cross-modal representation of the abstract notion of “several”
Figure 14: Text concept that is surface-level and visually-directed. Concept #2826 in SigLIP fires for text that describes blue objects. This is a surface-level text feature, describable by the identity of one token. It is also a feature of the text that relates directly to the visual modality.
Figure 15: Abstract linguistic text concept. Concept #1926 in SigLIP fires for text that contains uncertainty and/or modal verbs like “would” or “should”. This concept is picking up on abstract linguistic features about hypothetical. There is no clear relationship in the visual modality between these captions (unlike the “blue” text), and instead there is a language feature that connects them.
B.2 Illustration of Bridge Score
Appendix C Training Details for SAEs
We report the full training configuration used for the sparse autoencoders (SAEs) in this work. All SAEs were trained on frozen activations from the COCO dataset, consisting of 600,000 image and text embeddings per model. Each SAE was trained for 30 epochs, with a batch size of 1024, totaling approximately 18 million training examples. The encoder consists of a single linear projection without hidden layers, mapping input activations to a 4096-dimensional concept space. Unless specified, we used sparsity $k=5$, enforcing exactly five non-zero, non-negative coefficients per input.
Optimization was performed using the AdamW optimizer with a cosine learning rate schedule. The learning rate warmed up from $1\times 10^{-6}$ to a peak of $5\times 10^{-4}$, then decayed back to $1\times 10^{-6}$. We used weight decay of $1\times 10^{-5}$ and applied gradient clipping with a maximum global norm of 1.0. All model embeddings were $\ell_{2}$-normalized prior to encoding. For image inputs, the original images were resized to 256 pixels on the shorter side and center-cropped to $224\times 224$ before encoding by the VLMs. Text inputs were processed using the default tokenizer for each model.
We observed consistent convergence of the reconstruction loss around epoch 10, but trained for 30 epochs to ensure full stability of the learned dictionaries. We used the Overcompletehttps://github.com/KempnerInstitute/overcomplete toolbox for training.